### v1

@@ 5.1.1 Non-normalized event-study effects
Estimates of $\text{AVSQ}_{\ell}$ and pre-trends estimates are shown in Figure 2 below. Being exposed to a weakly larger number of newspapers for one electoral cycle increases turnout by 1.44 percentage point, and the effect is statistically significant (s.e.=0.43 percentage point). That effect can be estimated for 1,119 out of the 1,195 counties in the data: 34 counties never experience a change in their number of newspapers, and 42 counties that do experience a change cannot be matched with a not-yet-switcher with the same number of newspapers at baseline. Being exposed to a weakly larger number of newspapers for two, three, and four electoral cycle also significantly increases turnout. Effects increase with exposure length, but one cannot reject the null that all effects are equal (p-value=0.40). As $\ell$ increases, effects mechanically apply to fewer and fewer counties, but the effect after four electoral cycles still applies to 917 counties. Pre-trend estimates are small and individually and jointly insignificant. However, their confidence intervals are quite large. While the first pre-trend estimator applies to 906 of the 1,119 counties for which $\widehat{\text{AVSQ}}_{1}$ is estimated, the fourth pre-trend estimator only applies to 447 of the 917 counties for which $\widehat{\text{AVSQ}}_{4}$ is estimated. The confidence interval of $\widehat{\text{AVSQ}}_{-4}$ is already quite large, but that of $\widehat{\text{AVSQ}}_{-5}$ is substantially larger, so we have very little power to detect differential trends over more than five election cycles. This is why we only report four placebo and four event-study estimators.
Figure 2: Effects of being exposed to a weakly larger number of newspapers for $\ell$ periods on turnout
Note: This figure shows non-normalized DID estimates of the effect of being exposed to a weakly larger number of newspapers for $\ell$ periods on turnout, as well as pre-trends estimates, computed using the data of Gentzkow et al. (2011) and the did_multiplegt_dyn Stata command. Standard errors are clustered at the county level. 95% confidence intervals are in red.
@@ 5.1.2 Normalized event-study effects
Note: This figure shows normalized DID estimates of the effect of newspapers on turnout, as well as normalized pre-trends estimates, computed using the data of Gentzkow et al. (2011) and the did_multiplegt_dyn Stata command. Standard errors are clustered at the county level. 95% confidence intervals are in red.
5.1.3 Testing static effects
@@
Notes: “N. obs.” indicates the number of counties to which $\widehat{\text{AVSQ}}_{\ell}\!\!{}^{o}$ applies
@@ 5.2 Standard distributed-lag regressions
Next, we consider a distributed-lag TWFE estimator with just one lag. We obtain $\widehat{\beta}_{0}=-0.0008$ (s.e.=0.0014=0.0014) and $\widehat{\beta}_{1}=0.0050$ (s.e.=0.0015=0.0015). Hence, according to this regression, increasing the current number of newspapers insignificantly reduces turnout by 0.08 percentage points. On the other hand, increasing its first lag significantly increases turnout by 0.5 percentage points. These results may look surprising: the normalized event-study effects suggested the opposite, namely that the current number of newspapers would have a larger impact than the number of newspaper four years before. The tests above also do not reject the hypothesis that only the current number of newspapers affects turnout.
@@ 5.3 Heterogeneous distributed-lag regressions
Next, we turn to the estimation of an heterogeneous distributed-lag regressions, following Section 4.2 above. We consider $L=1$, as above. [^5] We also estimate the model with $L=2$ and $L=3$. However, already with $L=2$, the average in $\widehat{\overline{B}}$ applies to 116 counties only, from a total of 1,147 counties. The results are displayed in Table 2 below. One interesting aspect is that with $L=1$, the average in $\widehat{\overline{B}}$ applies to 1,071 counties, namely 93.3% of the sample. Nonetheless and perhaps not surprisingly, the estimates are somewhat imprecise, with standard errors more than twice as large as those of the usual TWFE distributed-lag regressions. With this caveat in mind, the results are much more in line with the event-study estimates above, and point towards a positive effect of the current number of newspapers, and little effects of the lagged number of newspapers.
We also estimate $\overline{B}(\mathcal{E})$ for $\mathcal{E}=\{D_{1}=0\}$, which holds for around 3/4 of the sample, and $\mathcal{E}=\{D_{1}>0\}$. The results point towards a larger effect of the current number newspaper when the county did not have any newspapers at the first period.
$\overline{B}$ $\overline{B}(\{D_{1}=0\})$ $\overline{B}(\{D_{1}>0\})$
$\ell=0$ $\ell=1$ $\ell=0$ $\ell=1$ $\ell=0$ $\ell=1$
Estimate 0.0043 -0.0009 0.0057 -0.0025 -0.0003 0.0043
Std. err. 0.0033 0.0034 0.0025 0.0026 0.0032 0.0029
N.obs 1071 822 249
Notes: the model is estimated on the 1,147 counties for which we have data on consecutive elections and no missing data on $D$. “N. obs” denotes the number of counties on which we average effects. Std. errors obtained by bootstrap.
Table 2: Estimates of $\overline{B}$, $\overline{B}(\{D_{1}=0\})$ and $\overline{B}(\{D_{1}>0\})$ with $L=1$
@@ A.3 Proof of Theorem 3
$\widehat{\beta}_{\ell}=\frac{\sum_{g=1}^{G}\sum_{t\geq L+1}\eta_{g,t}^{\ell}Y_{g,t}}{\sum_{g=1}^{G}\sum_{t\geq L+1}\eta_{g,t}^{\ell}D_{g,t-\ell}}.$
Moreover, by Assumption 7, for all $g$ and $t\geq L+1$,
$Y_{g,t}=Y_{g,t}(\bm{0}_{t})+\sum_{\ell=0}^{L}\beta_{g,\ell}D_{g,t-\ell}.$
By Assumptions 2 and 4 and the fact that $\eta_{g,t}^{\ell}$ is a residual, we have
$E\left[\frac{\sum_{g=1}^{G}\sum_{t\geq L+1}\eta_{g,t}^{\ell}Y_{g,t}(\bm{0}_{t})}{\sum_{g=1}^{G}\sum_{t\geq L+1}\eta_{g,t}^{\ell}D_{g,t-\ell}}\big{|}\bm{D}_{1},...,\bm{D}_{G}\right]$ $=\frac{\sum_{g=1}^{G}\sum_{t\geq L+1}\eta_{g,t}^{\ell}E[Y_{g,t}(\bm{0}_{t})|\bm{D}_{g}]}{\sum_{g=1}^{G}\sum_{t\geq L+1}\eta_{g,t}^{\ell}D_{g,t-\ell}}$
@@
Hence, by definition of $W_{g}^{\ell,\ell^{\prime}}$,
$E\left[\widehat{\beta}_{\ell}\big{|}\bm{D}_{1},...,\bm{D}_{G}\right]$ $=E\left[\sum_{\ell^{\prime}=0}^{L}\frac{\sum_{g=1}^{G}\sum_{t\geq L+1}\eta_{g,t}^{\ell}D_{g,t-\ell^{\prime}}\beta_{g,\ell^{\prime}}}{\sum_{g=1}^{G}\sum_{t\geq L+1}\eta_{g,t}^{\ell}D_{g,t-\ell}}\big{|}\bm{D}_{1},...,\bm{D}_{G}\right]$
$=\frac{1}{G}\sum_{g=1}^{G}E\left[\sum_{\ell^{\prime}=0}^{L}W_{g}^{\ell,\ell^{\prime}}\beta_{g,\ell}\big{|}\bm{D}_{1},...,\bm{D}_{G}\right].$
@@
$E\left[\widehat{\beta}_{\ell}\right$] $=\frac{1}{G}\sum_{g=1}^{G}E\left[\sum_{\ell^{\prime}=0}^{L}W_{g}^{\ell,\ell^{\prime}}\beta_{g,\ell}\right]$
$=E\left[\sum_{\ell^{\prime}=0}^{L}W^{\ell,\ell^{\prime}}\beta_{\ell}\right],$
where the last equality follows since groups are identically distributed. Moreover, remark that $(1/G)\sum_{g=1}^{G}W_{g}^{\ell,\ell}=1$. Since, again, groups are identically distributed, this implies that $E[W_{g}^{\ell,\ell}]=1$. Similarly, since $\eta_{g,t}^{\ell}$ is a residual, $(1/G)\sum_{g=1}^{G}W_{g}^{\ell,\ell^{\prime}}=0$ for all $\ell^{\prime}\neq\ell$, which implies that $E[W_{g}^{\ell,\ell^{\prime}}]=0$. Finally, if $\text{Cov}(W^{\ell,\ell^{\prime}},\beta_{\ell^{\prime}})=0$ for all $\ell^{\prime}\in\{0,...,L\}$,
$E\left[\widehat{\beta}_{\ell}\right$]= $\sum_{\ell^{\prime}=0}^{L}E\left[W^{\ell,\ell^{\prime}}\beta_{\ell^{\prime}}\right]$
$=$ $\sum_{\ell^{\prime}=0}^{L}E\left[W^{\ell,\ell^{\prime}}\right]E\left[\beta_{\ell^{\prime}}\right]$
$=$ $E\left[\beta_{\ell}\right].$
@@ A.4 Illustration of (17)
If $T=2$ and $L=0$, we simply have
$E\left[\|(\bm{M}\bm{M}^{\prime})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\|\,\big{|}\lambda_{\min}(\bm{M}\bm{M}^{\prime})>0\right]$
$=$ $E\left[\left|\frac{\Delta Y_{2}-\Gamma}{\Delta D_{2}}\right|\,\big{|}\Delta D_{2}\neq 0\right]$
$=$ $E\left[|B_{2}+\varepsilon_{2}/\Delta D_{2}|\,\big{|}\Delta D_{2}\neq 0\right]$
$\geq$ $E\left[\big{|}|B_{2}|-|\varepsilon_{2}/\Delta D_{2}|\big{|}\,|\Delta D_{2}\neq 0\right]$
$\geq$ $\big{|}E\left[|B_{2}|\,\big{|}\Delta D_{2}\neq 0\right]-E\left[|\varepsilon_{2}/\Delta D_{2}|\,\big{|}\Delta D_{2}\neq 0\right]\big{|}.$
Now, assume that $E[|\varepsilon_{2}|\,\big{|}\Delta D_{2}]\geq s>0$ a.s. and the distribution of $|\Delta D_{2}|$ is absolutely continuous with respect to the Lebesgue measure, with density $f$ satisfying $\lim_{x\downarrow 0}f(x)>0$. Then,
@@ A.5 Proof of Theorem 4
We need to show (18), and (16) when $D_{t}$ is finitely supported for all $t$. Regarding the first point, let us first show that
$E\left[\|(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\|\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>\lambda\right]<\infty.$ (20)
First, note that, letting $\|\cdot\|_{2}$ denote the spectral norm ($\|A\|_{2}=\lambda_{\max}(A^{\prime}A)^{1/2}$, with $\lambda_{\max}(A^{\prime}A)$ the largest eigenvalue of $A^{\prime}A$), we have
$\|(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\|=$ $\|(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\bm{M}B_{2}+\bm{\varepsilon})\|$
$\leq$ $\|B_{2}\|+\|(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}\bm{\varepsilon}\|$
$\leq$ $\|B_{2}\|+\|(\bm{M}^{\prime}\bm{M})^{-1}\|_{2}\times\|\bm{M}\|_{2}\times\|\bm{\varepsilon}\|$ (21)
Thus, conditional on $\lambda_{\min}(\bm{M}^{\prime}\bm{M})|>\lambda$, we have
$\|(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\|\leq\|B_{2}\|+\frac{1}{\lambda}\|\bm{M}\|_{2}\times\|\bm{\varepsilon}\|.$
Eq. (20) follows since $E[\|B_{2}\|]<\infty$ by Assumption 7, $E[D_{t}^{2}]<\infty$ for all $t\geq 1$, $E[\Delta Y_{t}(\bm{0}_{t})^{2}]<\infty$ for all $t\geq 2$ and Cauchy-Schwarz inequality.
Since (20) holds, $E\left[(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\,|\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>\lambda\right]$ is well-defined. Also,
$E\left[(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\,|\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>\lambda\right]=$ $E\left[B_{2}\,\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>\lambda\right]$
$+$ $E\left[(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}\Delta\varepsilon\,|\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>\lambda\right]$
$=$ $E\left[B_{2}\,\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>\lambda\right].$ (22)
Also, by the dominated convergence theorem,
$\lim_{\lambda\downarrow 0}E[B_{2}1\left\{\lambda_{\min}(\bm{M}^{\prime}\bm{M})|>\lambda\right\}]=$ $E[B_{2}1\left\{\lambda_{\min}(\bm{M}^{\prime}\bm{M})|>0\right\}],$
$\lim_{\lambda\downarrow 0}E[1\left\{\lambda_{\min}(\bm{M}^{\prime}\bm{M})|>\lambda\right\}]=$ $E[1\left\{\lambda_{\min}(\bm{M}^{\prime}\bm{M})|>0\right\}],$
which imply
$\lim_{\lambda\downarrow 0}E[B_{2}\lambda_{\min}(\bm{M}^{\prime}\bm{M})>\lambda]=\overline{B}.$
Equation (18) follows by combining the last display with (22).
Finally, we prove that (16) holds when $D_{t}$ is finitely supported for all $t$. It suffices to prove that in this case,
$E\left[\|(\bm{M}^{\prime}\bm{M})^{-1}\bm{M}^{\prime}(\Delta\bm{Y}-\Gamma)\|\,|\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0\right]<\infty.$ (23)
From (21), it suffices to show that
$E\left[\|(\bm{M}^{\prime}\bm{M})^{-1}\|_{2}\times\|\bm{M}\|_{2}\times\|\bm{\varepsilon}\|\,\big{|}\,\lambda_{\min}(\bm{M}^{\prime}\bm{M})>0\right]<\infty.$
Since $D_{t}$ is finitely supported for all $t$, $\|(\bm{M}^{\prime}\bm{M})^{-1}\|_{2}$ and $\|\bm{M}\|_{2}$ are also finitely supported, and therefore bounded. Eq. (23), and then (16), follow.

### v3

@@ 5.1.1 Non-normalized event-study effects
Estimates of $\text{AVSQ}_{\ell}$ and pre-trends estimates are shown in Figure 2 below. Being exposed to a weakly larger number of newspapers for one electoral cycle increases turnout by 1.44 percentage point, and the effect is statistically significant (s.e.=0.43 percentage point). That effect can be estimated for 1,119 out of the 1,195 counties in the data: 34 counties never experience a change in their number of newspapers, and 42 counties that do experience a change cannot be matched with a not-yet-switcher with the same number of newspapers at baseline. Being exposed to a weakly larger number of newspapers for two, three, and four electoral cycles also significantly increases turnout. Effects increase with exposure length, but one cannot reject the null that all effects are equal (p-value=0.40). As $\ell$ increases, effects mechanically apply to fewer and fewer counties, but the effect after four electoral cycles still applies to 917 counties.
Pre-trend estimates are small and individually and jointly insignificant. However, their confidence intervals are quite large. While the first pre-trend estimator applies to 906 of the 1,119 counties for which $\widehat{\text{AVSQ}}_{1}$ is estimated, the fourth pre-trend estimator only applies to 447 of the 917 counties for which $\widehat{\text{AVSQ}}_{4}$ is estimated. The confidence interval of $\widehat{\text{AVSQ}}_{-4}$ is already quite large, but that of $\widehat{\text{AVSQ}}_{-5}$ is substantially larger, so we have very little power to detect differential trends over more than five election cycles. This is why we only report four placebo and four event-study estimators.
Figure 2: Non-normalized DID estimates of effects of newspapers on turnout
Notes: we computed $\ell\mapsto\widehat{\text{AVSQ}}_{\ell}$ using the did_multiplegt_dyn Stata command. Vertical lines: 95% confidence intervals, based on standard errors clustered at the county level.
@@ 5.1.2 Normalized event-study effects
Notes: we computed $\ell\mapsto\widehat{\text{AVSQ}}^{n}_{\ell}$ using the did_multiplegt_dyn Stata command. Vertical lines: 95% confidence intervals, based on standard errors clustered at the county level.
5.1.3 Testing the null that lagged treatments do not affect the outcome
@@
Notes: “N. obs.” indicates the number of counties to which $\widehat{\text{AVSQ}}_{\ell}\!\!{}^{o}$ applies. Standard errors clustered at the county level.
@@ 5.2 Standard distributed-lag regressions
Next, we consider a distributed-lag TWFE estimator with just one lag. We obtain $\widehat{\beta}_{0}=-0.0008$ (s.e.=0.0014=0.0014) and $\widehat{\beta}_{1}=0.0050$ (s.e.=0.0015=0.0015). Hence, according to this regression, increasing the current number of newspapers insignificantly reduces turnout by 0.08 percentage points. On the other hand, increasing its first lag significantly increases turnout by 0.5 percentage points. These results may look surprising: the normalized event-study effects suggested the opposite, namely that the current number of newspapers would have a larger impact than the number of newspapers four years before. The tests above also do not reject the hypothesis that only the current number of newspapers affects turnout.
@@ 5.3 Heterogeneous distributed-lag regressions
Next, we turn to the estimation of heterogeneous distributed-lag regressions, following Section 4.2 above. We consider both $K=1$, as above, and $K=2$. The results are displayed in Table 2 below. Note that in both cases, the averages in $\widehat{\overline{\beta}}_{k}$ apply to a large part of the sample: even $\widehat{\overline{\beta}}_{2}$ applies to 1,037 counties, namely 90.4% of the sample. The estimators’ standard errors are around 50% larger than those of the distributed lag TWFE estimates, but they remain relatively precise. Finally, the results are much more in line with the event-study estimates above, and point towards a positive effect of the current number of newspapers, and little effects of the lagged number of newspapers.
Table 2: Estimates of the average effect of each lag
Model $K=1$ $K=2$
Parameter $\overline{\beta}_{0}$ $\overline{\beta}_{1}$ $\overline{\beta}_{0}$ $\overline{\beta}_{1}$ $\overline{\beta}_{2}$
Estimate 0.0045 -0.0008 0.0042 -0.0014 -0.0016
Std. err. 0.0021 0.0021 0.0021 0.0024 0.0022
N.obs 1,093 1,090 1,052 1,049 1,037
Notes: the two models are estimated on the 1,147 counties for which we have data on consecutive elections and no missing data on $D$. “N. obs” denotes the number of counties on which we average effects. Standard errors obtained by bootstrapping counties.
@@ A.3 Proof of Theorem 3
$\widehat{\beta}_{k}=\frac{\sum_{g=1}^{G}\sum_{t\geq K+1}\eta_{g,t}^{k}Y_{g,t}}{\sum_{g=1}^{G}\sum_{t\geq K+1}\eta_{g,t}^{k}D_{g,t-k}}.$
Moreover, by Assumption 7, for all $g$ and $t\geq K+1$,
$Y_{g,t}=Y_{g,t}(\bm{0}_{t})+\sum_{k=0}^{K}\beta_{g,k}D_{g,t-k}.$
By Assumptions 2 and 4 and the fact that $\eta_{g,t}^{k}$ is a residual, we have
$E\left[\frac{\sum_{g=1}^{G}\sum_{t\geq K+1}\eta_{g,t}^{k}Y_{g,t}(\bm{0}_{t})}{\sum_{g=1}^{G}\sum_{t\geq K+1}\eta_{g,t}^{k}D_{g,t-k}}\big|\bm{D}_{1},...,\bm{D}_{G}\right]$ $=\frac{\sum_{g=1}^{G}\sum_{t\geq K+1}\eta_{g,t}^{k}E[Y_{g,t}(\bm{0}_{t})|\bm{D}_{g}]}{\sum_{g=1}^{G}\sum_{t\geq K+1}\eta_{g,t}^{k}D_{g,t-k}}$
@@
Hence, by definition of $W_{g}^{k,k^{\prime}}$,
$E\left[\widehat{\beta}_{k}\big|\bm{D}_{1},...,\bm{D}_{G}\right]$ $=E\left[\sum_{k^{\prime}=0}^{K}\frac{\sum_{g=1}^{G}\sum_{t\geq K+1}\eta_{g,t}^{k}D_{g,t-k^{\prime}}\beta_{g,k^{\prime}}}{\sum_{g=1}^{G}\sum_{t\geq K+1}\eta_{g,t}^{k}D_{g,t-k}}\big|\bm{D}_{1},...,\bm{D}_{G}\right]$
$=\frac{1}{G}\sum_{g=1}^{G}E\left[\sum_{k^{\prime}=0}^{K}W_{g}^{k,k^{\prime}}\beta_{g,k}\big|\bm{D}_{1},...,\bm{D}_{G}\right].$
@@
$E\left[\widehat{\beta}_{k}\right]$ $=\frac{1}{G}\sum_{g=1}^{G}E\left[\sum_{k^{\prime}=0}^{K}W_{g}^{k,k^{\prime}}\beta_{g,k}\right]$
$=E\left[\sum_{k^{\prime}=0}^{K}W^{k,k^{\prime}}\beta_{k}\right],$
where the last equality follows since groups are identically distributed. Moreover, remark that $(1/G)\sum_{g=1}^{G}W_{g}^{k,k}=1$. Since, again, groups are identically distributed, this implies that $E[W_{g}^{k,k}]=1$. Similarly, since $\eta_{g,t}^{k}$ is a residual, $(1/G)\sum_{g=1}^{G}W_{g}^{k,k^{\prime}}=0$ for all $k^{\prime}\neq k$, which implies that $E[W_{g}^{k,k^{\prime}}]=0$. Finally, if $\text{Cov}(W^{k,k^{\prime}},\beta_{k^{\prime}})=0$ for all $k^{\prime}\in\{0,...,K\}$,
$E\left[\widehat{\beta}_{k}\right]=$ $\sum_{k^{\prime}=0}^{K}E\left[W^{k,k^{\prime}}\beta_{k^{\prime}}\right]$
$=$ $\sum_{k^{\prime}=0}^{K}E\left[W^{k,k^{\prime}}\right]E\left[\beta_{k^{\prime}}\right]$
$=$ $E\left[\beta_{k}\right].$
@@ A.4 Illustration of (17)
If $T=2$ and $K=0$, we simply have $\bm{M}=\bm{M}^{\prime}=[\Delta D_{2}]$ and $e_{0}\in\text{Im}(\bm{M}^{\prime})$ is equivalent to $\Delta D_{2}\neq 0$, in which case $f_{0}=1/\Delta D_{2}$. Then,
$E\left[\left|e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)\right|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\right]=$ $E\left[\left|\frac{\Delta Y_{2}-\gamma_{2}}{\Delta D_{2}}\right|\,\big|\Delta D_{2}\neq 0\right]$
$=$ $E\left[|\beta_{0}+\varepsilon_{2}/\Delta D_{2}|\,\big|\Delta D_{2}\neq 0\right]$
$\geq$ $E\left[\big||\beta_{0}|-|\varepsilon_{2}/\Delta D_{2}|\big|\,|\Delta D_{2}\neq 0\right]$
$\geq$ $\big|E\left[|\beta_{0}|\,\big|\Delta D_{2}\neq 0\right]-E\left[|\varepsilon_{2}/\Delta D_{2}|\,\big|\Delta D_{2}\neq 0\right]\big|.$
Now, if $E[|\varepsilon_{2}|\,\big|\Delta D_{2}]\geq s>0$ a.s. and the distribution of $|\Delta D_{2}|$ is absolutely continuous with respect to the Lebesgue measure, with density $f$ satisfying $\lim_{x\downarrow 0}f(x)>0$, then
@@ A.5 Proof of Theorem 4
We need to show (18), and (16) when $D_{t}$ is finitely supported for all $t$. Regarding the first point, first remark that by the dominated convergence theorem,
$\lim_{C\uparrow\infty}E\left[\beta_{k}1\left\{\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right\}|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime})\right]=$ $\overline{\beta}_{k},$ (23)
$\lim_{C\uparrow\infty}P\left(1\left\{\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right\}|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime})\right)=$ $1.$ (24)
This implies that for $C$ large enough, $P\left(e_{k}\in\text{Im}(\bm{M}^{\prime}),1\left\{\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right\}\right)>0$. We consider such $C$ below. Now, recall that if $e_{k}\in\text{Im}(\bm{M}^{\prime})$, then $\bm{M}^{\prime}\bm{M}^{\prime}{}^{+}e_{k}=e_{k}$. As a result, $e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)=\beta_{k}+e^{\prime}_{k}\bm{M}^{+}\bm{\varepsilon}$. Hence,
$E\left[\left|e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)\right|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right]$
$\leq$ $E\left[|\beta_{k}|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right]+ME\left[\|\bm{\varepsilon}\|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right]<\infty.$
Thus, for $C$ large enough, the expectation on the right-hand side of (18) is well-defined. Moreover,
$E\left[e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right]$ $=E\left[\beta_{k}\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right]$
$\hskip-11.38092pt+E\left[e^{\prime}_{k}\bm{M}^{+}\bm{\varepsilon},\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right]$
$=E\left[\beta_{k}\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C\right]$
$\to\overline{\beta}_{k},$
where the last line follows by (23)-(24).
Turning to the second point, remark that when $D_{t}$ is finitely supported, $\bm{M}^{\prime}$ and then $\bm{M}^{\prime}{}^{+}e_{k}$ are also finitely supported. Hence, $\|\bm{M}^{\prime}{}^{+}e_{k}\|$ is bounded with probability one and from what precedes,
$E\left[\left|e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)\right|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime})\right]=$ $E\left[\left|e^{\prime}_{k}\bm{M}^{+}(\Delta Y-\Gamma)\right|\,\big|\,e_{k}\in\text{Im}(\bm{M}^{\prime}),\|\bm{M}^{\prime}{}^{+}e_{k}\|\leq C_{0}\right]$
$<$ $\infty,$
for all $C$ large enough. Then, the right-hand side of (16) is well-defined, and the equality holds.